# MetaMath: Bootstrap Your Own Mathematical Questions for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have pushed the limits of natural language understanding and exhibited excellent problem-solving ability. Despite the great success, most existing open-source LLMs (e.g., LLaMA-2) are still far away from satisfactory for solving mathematical problem due to the complex reasoning procedures. To bridge this gap, we propose MetaMath, a fine-tuned language model that specializes in mathematical reasoning. Specifically, we start by bootstrapping mathematical questions by rewriting the question from multiple perspectives without extra knowledge, which results in a new dataset called MetaMathQA. Then we fine-tune the LLaMA-2 models on MetaMathQA. Experimental results on two popular benchmarks (i.e., GSM8K and MATH) for mathematical reasoning demonstrate that MetaMath outperforms a suite of open-source LLMs by a significant margin. Our MetaMath-7B model achieves $66.4\%$ on GSM8K and $19.4\%$ on MATH, exceeding the state-of-the-art models of the same size by $11.5\%$ and $8.7\%$. Particularly, MetaMath-70B achieves an accuracy of $82.3\%$ on GSM8K, slightly better than GPT-3.5-Turbo. We release the MetaMathQA dataset, the MetaMath models with different model sizes and the training code for public use.
*Equal contribution  Corresponding author
Figure 1: Overview of the MetaMathQA dataset and the mathematical problem-solving LLM – MetaMath. We note that our MetaMath-70B is fine-tuned by QLoRA [12] due to the compute limitation.
@@ 1 Introduction
Recent years have witnessed the rapid development of large language models (LLMs) which emerge as the favored approach for various applications in natural language processing, including open-ended conversational interfaces [5, 45, 44], coding assistance [6, 41, 35] and mathematical problem-solving [64, 24, 34, 11]. Among various tasks, solving mathematical problems is more challenging as they often require highly complex and symbolic multi-step reasoning capabilities. Although some close-sourced models, e.g., GPT-3.5-Turbo [42], GPT-4 [44] and PaLM-2 [57], have demonstrated promising performance on some mathematical problem-solving benchmarks, it is still a mystery how these models are trained and what data these models use. Therefore, how to equip open-source LLMs (e.g., LLaMA [56, 57]) with good mathematical problem-solving skills remains an open challenge.
@@
To tackle this challenge, two popular lines of research to improve the mathematical problem-solving abilities of LLMs are: prompt-based methods and finetuning-based methods. Prompt-based methods [62, 16, 61, 68, 16, 61] aim to activate the potential capacities of LLMs by choosing suitable prompting inputs without modifying the model parameters. Finetuning-based methods update the open-source LLMs (e.g., LLaMA) under the guidance of some other powerful closed-source LLMs (e.g., GPT-3.5 [42], GPT-4 [44]). While prompt-based methods are model-dependent and sensitive to many factors, finetuning-based methods, despite being simple and model-agnostic, heavily rely on effective training data on downstream mathematical questions. Our work aims to improve finetuning-based methods with a novel method to bootstrap available mathematical questions in the training set. Specifically, we propose to bootstrap the questions in both forward and backward reasoning directions. For the forward direction, we have the original and LLM-rephrased questions. For the backward direction, we have the self-verification and “IF” questions [63, 26]. To construct backward reasoning questions, we mask a token in a question using an identifier “x” and ask the model to predict the masked token if the answer is provided. Different from [63, 26] that apply backward reasoning for inference verification, we use it as a form of question for language model fine-tuning. For answers, we adopt an answer augmentation method based on rejection sampling [64], where diverse reasoning paths are generated and only those with correct answers are used. After combining both forward and backward mathematical questions with augmented answers, we construct a new dataset for fine-tuning, called MetaMathQA. By fine-tuning LLaMA-2 on MetaMathQA, we obtain our MetaMath model. Our approach is guided by the insight that a mathematical question represents merely a single view of the underlying meta-knowledge. Therefore, question bootstrapping can be viewed as a form of multi-view augmentation in order to enable the transfer of the meta-knowledge. Leveraging the MetaMathQA dataset, MetaMath demonstrates exceptional performance in mathematical reasoning, positioning it among the top performers on widely recognized evaluation benchmarks.
@@
Figure 2: GSM8K accuracy of LLaMA-2-7B fine-tuned on different sizes of answer augmentation data. Larger diversity gain indicates the question is more diverse compared to the existing questions. Detailed experimental setup is given in Section 4.1.
@@
Another motivation behind question bootstrapping is to enlarge the question diversity [14] such that the question distribution can be rich enough to cover more unseen scenarios. We quantify the question diversity of the original questions and our MetaMathQA dataset in Figure 2. The diversity gain [4] indicates how diverse the question is compared to the existing dataset, and larger diversity gain means the new question is more different from the existing dataset. With question bootstrapping, our MetaMathQA dataset is much more diverse than the original dataset. We also observe that the test accuracy without bootstrapped questions rapidly reaches a state of saturation. In contrast, the test accuracy, when using bootstrapped questions, continues to exhibit a steady increase.
@@
Question bootstrapping also has an intrinsic connection to dataset distillation [60, 66] and machine teaching [69, 32, 48], where the shared target is to construct a training dataset that best facilitates generalization. Unlike both methods that focus on optimizing the training empirical risk, question bootstrapping uses the reasoning diversity of questions as a heuristic proxy and maximizes this diversity by constructing forward, backward and rephrased questions. MetaMath aims to transfer the underlying meta-knowledge to enable strong generalization [28]. Our contributions are listed below:
@@
  • Based on the MetaMathQA dataset, MetaMath is fine-tuned from state-of-the-art open-source LLMs (e.g., LLaMA-2), showing excellent elementary mathematical problem-solving capability.
@@
  • We conduct experiments on two standard mathematical reasoning benchmarks: GSM8K [10] and MATH [19]. MetaMath outperforms existing open-source LLMs by a large margin. MetaMath-7B has achieved $66.4\%$ on GSM8K ($+11.5\%$ compared to the previous best open-source LLM) on GSM8K and $19.4\%$ on MATH ($+8.7\%$ compared to the previous best open-source LLM).
@@ 2 Related Work
Large Language Models (LLMs) [13, 33, 49, 50, 5, 56] have achieved great success in various natural language processing tasks, e.g., topic classification [38, 27], sentiment classification [5, 38], translation [5], by few-shot prompting (or in-context learning) [5, 38, 7]. Recently, Wei et al. [62], Wang et al. [61] show that LLMs with more than 100B parameters (e.g., GPT-3 [5] with 175B, PaLM with 540B [9]) can solve complex tasks by generating multiple reasoning steps towards the answer when given a few reasoning examples as demonstration. While both GPT-3.5 [42] and GPT-4 [44] have shown promising reasoning ability for complex mathematical tasks like MATH [19], the performance of open-source models (e.g., LLaMA-1 [56], LLaMA-2 [57]) is far from satisfactory.
@@
Learning Mathematical Reasoning for complex math tasks like GSM8K [10] and MATH [19] is one of the most challenging problem in open-source LLMs. Prompt-based methods aim to activate the potential capacities of LLMs by providing suitable prompting inputs without the need to perform model finetuning. Wei et al. [62] enhances the reasoning ability of LLMs by augmenting the output with a sequence of intermediate steps towards the answer. A few methods [16, 61, 68] are proposed to improve the quality of reasoning paths. For example, Complexity-based CoT [16] selects examples with more steps as in-context demonstrations and shows that prompting with more reasoning steps leads to better performance. Self-Consistency [61] samples multiple reasoning paths and selects the final answer by majority voting. Another category of works is finetuning-based methods, which fine-tunes open-source models (e.g., LLaMA) with the knowledge from some advanced closed-source LLMs [42, 44]. Magister et al. [36] investigates the transfer of reasoning capabilities via knowledge distillation. Yuan et al. [64] proposes to apply rejection sampling finetuning (RFT) to improve mathematical reasoning performance. WizardMath [34] proposes a reinforced evol-instruct method to enhance the reasoning abilities by supervised finetuning and PPO training [51]. Wang et al. [59] propose a constraint alignment loss to fine-tune LLMs for calibration.
@@
Knowledge Distillation [20, 17] transfers knowledge from a larger teacher model to a smaller student model, achieving promising performance in many applications [52, 46, 18, 39], Recently, [30, 23, 21, 36, 22, 15, 53] propose to transfer reasoning abilities from LLMs (e.g., GPT-3.5 [42], PaLM [9]) to small language models (e.g., T5 [50], GPT-2 [49]). For example, Finetune-CoT [21] samples multiple reasoning paths from LLMs and finetune student models with correct ones, while Self-Improve [23] chooses the one with the highest confidence. Li et al. [30] further feeds the question and ground-truth label to LLMs for prompting its reasoning path. Shridhar et al. [53] proposes to generate sub-questions and solution pairs for training. Small models finetuned by knowledge distillation can achieve a similar performance to LLMs [36, 21] on both common sense reasoning (e.g., CommonSenseQA [54]) and symbol reasoning (e.g., Coin Flip [62]). However, for solving challenging mathematical problems (e.g., GSM8K [10]), there is still a large performance gap [21, 15, 36].
@@ 3 Method
The overview of our method is illustrated in Figure 1. Given a meta-question (a sample in the original mathematical training set), we can generate a series of variants. Specifically, we perform three types of question bootstrapping. Combined with answer augmentation, we present MetaMathQA, a diverse and high-quality mathematical dataset based on GSM8K and MATH. We then present MetaMath, a familty of LLMs fine-tuned on MetaMathQA focusing on elementary mathematical problem-solving.
@@ 3.1 Answer Augmentation (AnsAug)
Generating more reasoning paths is a simple but effective way to augment the training set. For a question $q_{i}$, we use few-shot chain-of-thought prompting with temperature sampling  [^1] Temperature is set to 0.7 as in [61]. to generate $K_{\text{AugAng}}$ more reasoning paths $\{(r_{i}^{(j)},a_{i}^{(j)}):j=1,\dots,K_{\text{AugAng}}\}$: the question is appended to a few in-context reasoning examples, then fed to the LLM for generating its reasoning path $r_{i}^{(j)}$ and answer $a_{i}^{(j)}$. We filter out reasoning paths with correct answers as:
$\mathcal{D}_{\text{AnsAug}}=\{(q_{i},r_{i}^{(j)},a_{i}^{(j)}):a_{i}^{(j)}=a_{i}^{\star};i=1,\dots,N_{q};j=1,\dots,K_{\text{AugAng}}\}.$ (1)
@@ 3.2 Question Bootstrapping by LLM Rephrasing
Generating more answers for mathematical questions with LLMs is straightforward, but creating questions is more challenging. The questions in GSM8K and MATH are written by well-educated teachers. Hence, enlarging the question set through manual creation is time-consuming and labor-intensive. To address this issue, we propose rephrasing prompting to generate more questions through the LLM. Example LABEL:exmp:rephrase-prompt (see Example in Appendix A.1 for the complete prompt) shows a few-shot prompt used in GSM8K. Specifically, for a question $q_{i}$, we append it to the prompt, which is then fed to the LLM for generating the rephrased question. We adopt temperature sampling to sample $K_{\text{rephrase}}$ rephrased questions for each meta-question. For the rephrased questions, it is time-consuming to manually check the consistency compared with the original questions. We propose a supervised method to evaluate the correctness between the rephrased questions and the meta-questions. For each rephrased question $\hat{q}_{i}^{(j)}$, we use few-shot Chain-of-Thought prompting to generate its reasoning path $\hat{r}_{i}^{(j)}$ and answer $\hat{a}_{i}^{(j)}$, which is compared with the ground-truth answer $a_{i}^{\star}$. The accuracy of Complexity-based CoT [16] for answering the rephrased question by GPT-3.5-Turbo is $76.30\%$, which is comparable to that of answering the original training questions ($80.74\%$). This suggests that the quality of rephrased questions is preserved high while the question diversity is improved. We collect the rephrased questions with correct answers (i.e., $\hat{a}_{i}^{(j)}=a_{i}^{\star}$) as the augmented data:
@@
{exmp}
Prompt for Rephrasing GSM8K Questionsrephrase-promptYou are an AI assistant to help me rephrase questions. Follow the given examples.
Question: Olivia has \$23. She bought five bagels for \$3 each. How much money does she have left?
Rephrase the above question: What is the amount of money that Olivia has left after purchasing five bagels for \$3 each, if she initially had \$23?
Question: Michael had 58 golf balls. On Tuesday, he lost 23 golf balls. On Wednesday, he lost 2 more. How many golf balls did he have at the end of Wednesday?
Rephrase the above question: After losing 23 golf balls on Tuesday and an additional 2 on Wednesday, how many golf balls does Michael have left if he initially had 58 golf balls? … Question: {Q}
Rephrase the above question:
@@ 3.3 Question Bootstrapping by Backward Reasoning
Backward reasoning plays an important role in answering many mathematical questions, i.e., starting with a given condition and thinking backward to determine an unknown variable in the question. Usually, these questions contain a “IF” condition statement, as shown in Example LABEL:exmp:if. In total, 35% of testing questions in GSM8K contain the “IF” statement. However, existing methods (SFT, RFT, WizardMath) have a significant lower accuracy on testing questions with “IF”, as shown in Figure 7.
{exmp}
Question with “IF”if Question: Mark has \$50 in his bank account. He earns \$10 per day at his work. If he wants to buy a bike that costs \$300, how many days does Mark have to save his money? Answer: Mike has \$50 in his bank account, so the amount he needs to save to reach the bike goal is \$300 - \$50 = \$250. He earns \$10 per day, so \$250 / \$10/day = <<much-less-than$<$ <250/10=25>>much-greater-than$>$ >25 days #### 25
To improve the backward reasoning ability of fine-tuned models, we generate more questions which can be solved in a backward manner: a number in the question $q_{i}$ is masked by “x”, while the LLM is asked to predict the value of “x” when its answer $a_{i}^{\star}$ is provided. Different from forward reasoning, which generates explicit intermediate steps towards the final answer, backward reasoning starts with the answer and generates multiple reasoning steps to predict the masked number. Representative backward reasoning methods include Self-Verification [63] and FOBAR [26].
In Self-Verification (SV) [63], the question with the answer is first rewritten into a declarative statement, e.g., “How much did she earn?” (with the answer 10) is rewritten into “She earned \$10”. Different from Self-Verification that rewrites the question in a zero-shot manner by an instruction prompt “Please change the questions and answers into complete declarative sentences {Q} The answer is:”, we propose a few-shot prompting to rewrite the questions. Specifically, we design a prompt with several in-context rewriting examples (question with answer, its declarative statement); then a new question with answer is appended to the prompt before feeding to the LLM for generating its declarative statement. Example LABEL:exmp:sv shows the template used for rewriting questions. Then, a question for asking the value of ${\bf x}$ is appended, e.g., “What is the value of unknown variable ${\bf x}$?”. Example LABEL:exmp:sv-example gives an augmented example. We collect the new questions and their generated reasoning paths with correct answers as the augmented data:
@@
{exmp}
Prompts for Rewriting Question with Answer into a Declarative Statementsv You are an AI assistant to help me rewrite question into a declarative statement when its answer is provided. Follow the given examples and rewrite the question.
Question: How many cars are in the parking lot? The answer is: 5.
Result: There are 5 cars in the parking lot.
Question: {Q} The answer is: {A}.
Result: {exmp}Self-Verification [63] Questionsv-example Question: James buys x packs of beef that are 4 pounds each. The price of beef is \$5.50 per pound. He paid 110. What is the value of unknown variable x?
Answer: To solve this problem, we need to determine the value of x, which represents the number of packs of beef that James bought. Each pack of beef weighs 4 pounds and costs \$5.50 per pound. The total amount James paid is \$110. We can set up the equation as follows: Number of packs of beef * Weight per pack * Price per pound = Total amount paid; x * 4 * \$5.50 = \$110; Let’s simplify and solve for x: 4x * \$5.50 = \$110; 22x = \$110; To isolate x, we divide both sides of the equation by 22: 22x / 22 = \$110 / 22; x = \$5; The value of x is 5. Self-Verification needs to rewrite the question with answer into a declarative statement, which is challenging for complex questions. To address this issue, FOBAR [26] proposes to directly append the answer to the question, i.e., “If we know the answer to the above question is {$a_{i}^{\star}$} , what is the value of unknown variable x?” Example LABEL:exmp:fobar shows an example. We collect the new questions along with their correct answers as our augmented data:
@@
{exmp}
FOBAR [26] QuestionfobarQuestion: James buys x packs of beef that are 4 pounds each. The price of beef is \$5.50 per pound. How much did he pay? If we know the answer to the above question is 110, what is the value of unknown variable x?
Answer: James buys x packs of beef that are 4 pounds each, so he buys a total of 4x pounds of beef. The price of beef is \$5.50 per pound, so the total cost of the beef is 5.50 * 4x = 22x. We are given that the total cost is \$110, so we can write: 22x = 110. Dividing both sides by 22, we get: x = 5. The value of x is 5.

### v4

@@ Abstract
Large language models (LLMs) have pushed the limits of natural language understanding and exhibited excellent problem-solving ability. Despite the great success, most existing open-source LLMs (e.g., LLaMA-2) are still far away from satisfactory for solving mathematical problems due to the complex reasoning procedures. To bridge this gap, we propose MetaMath, a finetuned language model that specializes in mathematical reasoning. Specifically, we start by bootstrapping mathematical questions by rewriting the question from multiple perspectives, which results in a new dataset called MetaMathQA. Then we finetune the LLaMA-2 models on MetaMathQA. Experimental results on two popular benchmarks (i.e., GSM8K and MATH) for mathematical reasoning demonstrate that MetaMath outperforms a suite of open-source LLMs by a significant margin. Our MetaMath-7B model achieves $66.5\%$ on GSM8K and $19.8\%$ on MATH, exceeding the state-of-the-art models of the same size by $11.5\%$ and $8.7\%$. Particularly, MetaMath-70B achieves an accuracy of $82.3\%$ on GSM8K, slightly better than GPT-3.5-Turbo. We release the MetaMathQA dataset, the MetaMath models with different model sizes and the training code for public use.
[…]
Project page: meta-math.github.io
Figure 1: Overview of the MetaMathQA dataset and the mathematical problem-solving LLM – MetaMath. We note that our MetaMath-70B is finetuned by QLoRA [15] due to the computing resource limitation.
@@ 1 Introduction
Recent years have witnessed the rapid development of large language models (LLMs) which emerge as the favored approach for various applications and demonstrate multi-dimensional abilities, including instruction following [7, 67, 55, 38], coding assistance [8, 51, 44, 36], and mathematical problem-solving [79, 28, 43, 14]. Among various tasks, solving mathematical problems is more challenging as they often require highly complex and symbolic multi-step reasoning capabilities. Although some close-sourced models, e.g., GPT-3.5-Turbo [52], GPT-4 [54] and PaLM-2 [70], have demonstrated promising performance on some mathematical problem-solving benchmarks, it is still a mystery how these models are trained and what data these models use. Therefore, how to equip open-source LLMs (e.g., LLaMA [69, 70]) with good mathematical problem-solving skills remains an open challenge.
@@
To tackle this challenge, two popular lines of research to improve the mathematical problem-solving abilities of LLMs are: prompt-based methods and finetuning-based methods. Prompt-based methods [75, 20, 74, 84, 78, 77] aim to activate the potential capacities of LLMs by choosing suitable prompting inputs without modifying the model parameters. Finetuning-based methods update the open-source LLMs (e.g., LLaMA) under the guidance of some other powerful closed-source LLMs (e.g., GPT-3.5 [52], GPT-4 [54]). While prompt-based methods are model-dependent and sensitive to many factors, finetuning-based methods, despite being simple and model-agnostic, heavily rely on effective training data on downstream mathematical questions. Our work aims to improve finetuning-based methods with a novel method to bootstrap available mathematical questions in the training set. Specifically, we propose to bootstrap the questions in both forward and backward reasoning directions. For the forward direction, we have the original and LLM-rephrased questions. For the backward direction, we have the self-verification question [76] and FOBAR question [32]. To construct backward reasoning questions, we mask a token in a question using an identifier “x” and ask the model to predict the masked token if the answer is provided. Different from [76, 32] that apply backward reasoning for inference verification, we use it as a form of question for language model fine-tuning. For answers, we adopt an answer augmentation method based on rejection sampling [79], where diverse reasoning paths are generated and only those with correct answers are used. After combining both forward and backward mathematical questions with augmented answers, we construct a new dataset for fine-tuning, called MetaMathQA. By fine-tuning LLaMA-2 on MetaMathQA, we obtain our MetaMath model. Our approach is guided by the insight that a mathematical question represents merely a single view of the underlying meta-knowledge. Therefore, question bootstrapping can be viewed as a form of multi-view augmentation in order to enable the transfer of the meta-knowledge. Leveraging the MetaMathQA dataset, MetaMath demonstrates exceptional performance in mathematical reasoning, positioning it among the top performers on widely recognized evaluation benchmarks.
@@
Figure 2: GSM8K accuracy of LLaMA-2-7B finetuned on different sizes of answer augmentation data. A larger diversity gain indicates the question is more diverse compared to the existing questions. Detailed experimental setup is given in Section 4.1.
@@
Another motivation behind question bootstrapping is to enlarge the question diversity [18] such that the question distribution can be rich enough to cover more unseen scenarios. We quantify the question diversity of the original questions and our MetaMathQA dataset in Figure 2. The diversity gain [6] indicates how diverse the question is compared to the existing dataset, and a larger diversity gain means the new question is more different from the existing dataset. With question bootstrapping, our MetaMathQA dataset is much more diverse than the original dataset. We also observe that the test accuracy without bootstrapped questions rapidly reaches a state of saturation. In contrast, the test accuracy, when using bootstrapped questions, continues to exhibit a steady increase.
@@
Question bootstrapping also has an intrinsic connection to dataset distillation [73, 82] and machine teaching [85, 40, 41, 58], where the shared target is to construct a training dataset that best facilitates generalization. Unlike both methods that focus on optimizing the training empirical risk, question bootstrapping uses the reasoning diversity of questions as a heuristic proxy and maximizes this diversity by constructing forward, backward and rephrased questions. MetaMath aims to transfer the underlying meta-knowledge to enable strong generalization [34]. Our contributions are listed below:
@@
  • Based on the MetaMathQA dataset, MetaMath is finetuned from state-of-the-art open-source LLMs (e.g., LLaMA-2), showing excellent elementary mathematical problem-solving capability.
@@
  • We conduct experiments on two standard mathematical reasoning benchmarks: GSM8K [13] and MATH [23]. MetaMath outperforms existing open-source LLMs by a large margin. MetaMath-7B has achieved $66.5\%$ on GSM8K ($+11.5\%$ compared to the previous best open-source LLM) on GSM8K and $19.8\%$ on MATH ($+8.7\%$ compared to the previous best open-source LLM).
@@ 2 Related Work
Large Language Models (LLMs) [16, 42, 59, 60, 7, 69, 65] have achieved great success in various natural language processing tasks, e.g., topic classification [47, 33, 31], sentiment classification [7, 47], translation [7], by few-shot prompting (or in-context learning) [7, 47, 10]. Recently, Wei et al. [75], Wang et al. [74] show that LLMs with more than 100B parameters (e.g., GPT-3 [7] with 175B, PaLM with 540B [12]) can solve complex tasks by generating multiple reasoning steps towards the answer when given a few reasoning examples as demonstration. While both GPT-3.5 [52] and GPT-4 [54] have shown promising reasoning ability for complex mathematical tasks like MATH [23], the performance of open-source models (e.g., LLaMA-1 [69], LLaMA-2 [70]) is far from satisfactory.
@@
Learning Mathematical Reasoning for complex math tasks like GSM8K [13] and MATH [23] is one of the most challenging problem in open-source LLMs. Wei et al. [75] enhances the reasoning ability of LLMs by augmenting the output with a sequence of intermediate steps toward the answer. A few methods [20, 74, 84] are proposed to improve the quality of reasoning paths. For example, Complexity-based CoT [20] selects examples with more steps as in-context demonstrations and shows that prompting with more reasoning steps leads to better performance. Self-Consistency [74] samples multiple reasoning paths and selects the final answer by majority voting. Another category of work is finetuning-based methods, which finetunes open-source models (e.g., LLaMA) with the knowledge from some advanced closed-source LLMs [52, 54]. Magister et al. [45] investigates the transfer of reasoning capabilities via knowledge distillation. Yuan et al. [79] proposes to apply rejection sampling finetuning (RFT) to improve mathematical reasoning performance. WizardMath [43] proposes a reinforced evol-instruct method to enhance reasoning abilities by supervised fine-tuning and PPO training [62]. MAmmoTH [80] combines CoT and Program-of-Thought [9] rationales for teaching LLMs to use external tools (e.g., Python interpreter) for solving mathematical problems. Wang et al. [72] propose a constraint alignment loss to finetune LLMs for calibration.
@@
Knowledge Distillation [24, 21] transfers knowledge from a larger teacher model to a smaller student model, achieving promising performance in many applications [63, 56, 22, 48], Recently, [37, 27, 25, 45, 26, 19, 64] propose to transfer reasoning abilities from LLMs (e.g., GPT-3.5 [52], PaLM [12]) to small language models (e.g., T5 [60], GPT-2 [59]). For example, Finetune-CoT [25] samples multiple reasoning paths from LLMs and finetune the student model with correct ones, while Self-Improve [27] chooses the one with the highest confidence. Li et al. [37] further feeds the question and ground-truth label to LLMs for prompting its reasoning path. Shridhar et al. [64] proposes to generate sub-questions and solution pairs for training. Small models finetuned by knowledge distillation can achieve similar performance to LLMs [45, 25] on both common sense reasoning (e.g., CommonSenseQA [66]) and symbol reasoning (e.g., Coin Flip [75]). However, for solving challenging mathematical problems (e.g., GSM8K [13]), there is still a large performance gap [25, 19, 45].
@@ 3 Method
The overview of our method is illustrated in Figure 1. Given a meta-question (a sample in the original mathematical training set), we can generate a series of variants. Specifically, we perform three types of question bootstrapping. Combined with answer augmentation, we present MetaMathQA, a diverse and high-quality mathematical dataset based on GSM8K and MATH. We then present MetaMath, a family of LLMs finetuned on MetaMathQA focusing on elementary mathematical problem-solving.
@@ 3.1 Answer Augmentation (AnsAug)
Generating more reasoning paths is a simple but effective way to augment the training set. For a question $q_{i}$, we use few-shot chain-of-thought prompting with temperature sampling to generate $K_{\text{AnsAug}}$ more reasoning paths $\{(r_{i}^{(j)},a_{i}^{(j)}):j=1,\dots,K_{\text{AnsAug}}\}$: the question is appended to a few in-context reasoning examples, then fed to the LLM for generating its reasoning path $r_{i}^{(j)}$ and answer $a_{i}^{(j)}$. We filter out reasoning paths with correct answers as:
$\mathcal{D}_{\text{AnsAug}}=\{(q_{i},r_{i}^{(j)},a_{i}^{(j)}):a_{i}^{(j)}=a_{i}^{\star};i=1,\dots,N_{q};j=1,\dots,K_{\text{AnsAug}}\}.$ (1)
@@ 3.2 Question Bootstrapping by LLM Rephrasing
Generating more answers for mathematical questions with LLMs is straightforward, but creating questions is more challenging. Math Questions are written by well-educated teachers. Hence, enlarging the question set through manual creation is time-consuming and labor-intensive. To address this issue, we propose rephrasing prompting to generate more questions through the LLM.
Example 3.1:  Rephrasing Question Question: What is the total amount that James paid when he purchased 5 packs of beef, each weighing 4 pounds, at a price of \$5.50 per pound? Answer: Each pack of beef weighs 4 pounds, so 5 packs weigh 4 * 5 = 20 pounds in total. The price per pound of beef is \$5.50, so the total cost for 20 pounds is 20 * \$5.50 = \$110.  …  The answer is: 110.
Specifically, for a question $q_{i}$, we append it to the prompt, which is then fed to the LLM for generating the rephrased question. Example 3.2 shows a generated rephrased question and the complete prompt is shown in Appendix A.1. We adopt temperature sampling to sample $K_{\text{rephrase}}$ rephrased questions for each meta-question. For the rephrased questions, it is time-consuming to manually check the consistency compared with the original questions. We propose a supervised method to evaluate the correctness between the rephrased questions and the meta-questions. For each rephrased question $\hat{q}_{i}^{(j)}$, we use few-shot Chain-of-Thought prompting to generate its reasoning path $\hat{r}_{i}^{(j)}$ and answer $\hat{a}_{i}^{(j)}$, which is compared with the ground-truth answer $a_{i}^{\star}$. The accuracy of Complexity-based CoT [20] for answering the rephrased question by GPT-3.5-Turbo is $76.30\%$, which is comparable to that of answering the original training questions ($80.74\%$). This suggests that the quality of rephrased questions is preserved high while the question diversity is improved. We collect the rephrased questions with correct answers (i.e., $\hat{a}_{i}^{(j)}=a_{i}^{\star}$) as the augmented data:
@@ 3.3 Question Bootstrapping by Backward Reasoning
Backward reasoning plays an important role in answering many mathematical questions, i.e., starting with a given condition and thinking backward to determine an unknown variable in the question. One specific example between a question and a backward question is illustrated in Example 3.3. However, existing methods (SFT, RFT, WizardMath) have significantly lower accuracy on backward questions, as shown in Figure 7, motivating us to bootstrap backward questions to improve the reasoning ability.
Example 3.2:  Question and Backward Question Question: James buys 5 packs of beef that are 4 pounds each. The price of beef is \$5.50 per pound. How much did he pay? Answer: He bought 5*4=20 pounds of beef. He paid 20*5.5=\$110. The answer is: 110 ✓ Backward Question: James buys x packs of beef that are 4 pounds each. The price of beef is \$5.50 per pound. How much did he pay? If we know the answer to the above question is 110, what is the value of unknown variable x? Answer: The total weight of the beef is 4*x because 4*5.5 = 22. … The answer is: 27 ✗
To improve the backward reasoning ability of finetuned models, we generate more questions which can be solved in a backward manner: a number in the question $q_{i}$ is masked by “x”, while the LLM is asked to predict the value of “x” when its answer $a_{i}^{\star}$ is provided. Different from forward reasoning, which generates explicit intermediate steps towards the final answer, backward reasoning starts with the answer and generates multiple reasoning steps to predict the masked number. Representative backward reasoning methods include Self-Verification [76] and FOBAR [32].
In Self-Verification (SV) [76], the question with the answer is first rewritten into a declarative statement, e.g., “How much did he pay?” (with the answer 110) is rewritten into “He paid \$10”. Then, a question for asking the value of ${\bf x}$ is appended, e.g., “What is the value of unknown variable ${\bf x}$?”. Example 3.3 gives an augmented example. We collect the new questions and their generated reasoning paths with correct answers as the augmented data:
@@
Example 3.3:  Self-Verification [76] Question Question: James buys x packs of beef that are 4 pounds each. The price of beef is \$5.50 per pound. He paid 110. What is the value of unknown variable x? Answer: To solve this problem, we need to determine the value of x, which represents the number of packs of beef that James bought. Each pack of beef weighs 4 pounds and  …  The value of x is 5.
Example 3.4:  FOBAR [32] Question Question: James buys x packs of beef that are 4 pounds each. The price of beef is \$5.50 per pound. How much did he pay? If we know the answer to the above question is 110, what is the value of unknown variable x? Answer: James buys x packs of beef that are 4 pounds each, so he buys a total of 4x pounds of beef. The price of beef is \$5.50 per pound, so the total cost of the beef is 5.50 * 4x = 22x.  …  The value of x is 5.
Self-Verification needs to rewrite the question with an answer into a declarative statement, which is challenging for complex questions. To address this issue, FOBAR [32] proposes to directly append the answer to the question, i.e., “If we know the answer to the above question is {$a_{i}^{\star}$} , what is the value of unknown variable x?” Example 3.3 shows an example. We collect the new questions along with their correct answers as our augmented data: